OpenStudy (anonymous):

g(1) = 4 g(n+1)=4[g(n-1)] Find a closed form formula.

OpenStudy (freckles):

have you tried to plug in other values of n and see if there is a pattern this is one way to go about the problem

OpenStudy (anonymous):

Yes, but the thing is when you plug it in, you have to know different y-values that are not given, and I do not know how to get those y-values.

OpenStudy (freckles):

so we know g(1)=4 n-1=1 when n=2 so choose n=2 we have g(2+1)=4g(2-1) g(3)=4g(1)=4*4=16 so right now we have g(1)=4 g(3)=16 ...we can find g(5) next by inserting n=4 ...we can find g(7) by then inserting n=6

OpenStudy (freckles):

now we got to figure out how to evaluate g(even integer)

OpenStudy (anonymous):

oh dang that makes sense.

OpenStudy (freckles):

I'm having trouble with g(even integer)

OpenStudy (anonymous):

Same.

OpenStudy (freckles):

@zepdrix what do you think? how to find g(2)?

OpenStudy (freckles):

can i ask what other techniques you have learned for finding closed forms?

OpenStudy (freckles):

if n-1 is odd then n+1 is odd and we are given g(1) which is g(odd number) if n-1 is even then n+1 is even but we are not given a condition for g(even number)=something so I feel like we only have enough information for a closed form for g(odd integer)

OpenStudy (anonymous):

Our teacher really did not say anything about finding them other than using the trend of the output

OpenStudy (freckles):

I can get to: \[g(2n+1)=4^{n+1} \text{ this is a closed from for the odd integers }\] hmm... we can try to play with this maybe...

OpenStudy (freckles):

we can assume this will hold for the even integers (this is may not be true though) x=2n+1 then n=(x-1)/2 So we have \[g(x)=4^{\frac{x-1}{2}+1}=4^{\frac{x-1+2}{2}}=4^{\frac{x+1}{2}} \\ g(1)=4^{\frac{1+1}{2}}=4^{1}=4 \\ g(2)=4^{\frac{2+1}{2}}=4^{\frac{3}{2}} \\ g(3)=4^\frac{3+1}{2}=4^\frac{4}{2}=4^2=16 \\ g(4)=4^{\frac{4+1}{2}}=4^\frac{5}{2}\] ...

OpenStudy (freckles):

which does make the conditions g(n+1)=4g(n-1) and g(1)=4 hold

OpenStudy (freckles):

you can verify thing by showing g(1)=4 using g(x)=4^((x+1)/2) and you can also show g(n+1)=4g(n-1) \[g(n+1)=4^\frac{(n+1)+1}{2} \\ g(n-1)=4^\frac{(n-1)+1}{2} \\ \text{ now inputting into } \\ g(n+1)=4g(n-1) \\ 4^\frac{(n+1)+1}{2}=4 \cdot 4^\frac{(n-1)+1}{2} \\ 4^\frac{n+2}{2}=4 \cdot 4^\frac{n}{2} \\ 4^{\frac{n+2}{2}}=4^{\frac{2}{2} } 4^{\frac{n}{2}} \\ 4^\frac{n+2}{2}=4^{\frac{2+n}{2}} \\ 4^{\frac{n+2}{2}}=4^{\frac{n+2}{2}} \text{ this is a true equation }\] ...so this form does work... but it is not the only form that works for what was given since we were are given any conditions on the even numbers

OpenStudy (freckles):

we were not given *